what is the answer to 5x+2=x+7
Answer by bucky(2097)

The general steps to solving an equation such as this are:
.
First collect all the terms containing x on one side of the equation and all the constants on
the other side.
.
Second combine terms to simplify the equation.
.
Third divide both sides by the multiplier of x to solve for x
.
Let's apply those steps.
.
Start with the given equation:
.
5x + 2 = x + 7
.
Get rid of the +2 on the left side by subtracting 2 from both sides. When you do that subtraction
the equation becomes:
.
5x + 2 - 2 = x + 7 - 2
.
Next get rid of the x on the right side by subtracting x from both sides. This makes the
equation become:
.
5x - x + 2 - 2 = x - x + 7 - 2
.
Now we simplify the results. On the left side the 5x - x becomes 4x and the +2 - 2 cancel each
other out making the 2's disappear. This first simplification reduces the left side, making the
equation:
.
4x = x - x + 7 -2
.
On the right side combining x and -x results in the x's disappearing on the right side.
And combining +7 and -2 results in +5. So now the equation is reduced to:
.
4x = 5
.
Finally, solve for x by dividing both sides by 4 ... because the multiplier of x is 4. This
division makes the equation become:
.
x = 5/4
.
That's the answer to this problem ... x = 5/4.
.
Hope this explanation helps you to understand the problem. When you get a little more practice
in working these problems, you will be able to do a lot of the detail in your head without
writing every step as was done above. That will make things a lot easier.
.